What is the line of symmetry for the graph of y=x^2-4x-5?
Answer by stanbon(75887)

What is the line of symmetry for the graph of y=x^2-4x-5?
-----------
Rewrite as:
x^2-4x+?=-y+5+?
Complete the square to get:
x^2-4x+4 = -y+5+4
(x+2)^2=-(y-9)
Axis of symmetry is x=-2
Cheers,
Stan H.